One positive number is 3 times another number. The difference between the two numbers is 62. what are the numbers?
Answer by macston(5194)

One is three times the another A=3B ;the difference is 62 A-B=62
3B-B=62 substitute first equation for A in second equation and solve for B
2B=62
B=31 ANSWER 1 B=31
A=3B
A=3(31)
A=93 ANSWER 2 A=93
CHECK:
A-B=62
93-31=62
62=62